Exercise 14.54 with feedback

For the following reaction, Kc = 255 at 1000 K.
CO (g) + Cl2 (g) ⇌ COCl2 (g)
A reaction mixture initially contains a COconcentration of 0.1550 M and a Cl2concentration of 0.171 M at 1000 K.

Part A

What is the equilibrium concentration of CO at 1000 K?

Express your answer in molarity to three significant figures.

[CO] =   M  

SubmitMy AnswersGive Up

Part B

What is the equilibrium concentration of Cl2 at 1000 K?

Express your answer in molarity to three significant figures.

[Cl2] =   M  

SubmitMy AnswersGive Up

Part C

What is the equilibrium concentration of COCl2 at 1000 K?

Express your answer in molarity to three significant figures.

[COCl2] =   M  

SubmitMy AnswersGive Up
Answer #1

I 0 IssO ML 017I E (0-1550-9 (0、M-X toeo k Given col = 0.13837- AL Part A r swi = 10.01661 M once ast & ct, こ (0、171-0 . 1383